OpenStudy (anonymous):

A website recorded 1,351 visitors one month. This was 30% less than the number of visitors to the site the previous month. The following equation represents this situation, where x represents the number of visitors to the site the previous month. x - 0.3x = 1,351 How many visitors did the website record in both months combined?

OpenStudy (australopithecus):

x - 0.3x = 1,351 x(1-0.3) = 1,351 x = 1,351/0.7

OpenStudy (australopithecus):

you should be able to solve from here

OpenStudy (anonymous):

3,281

OpenStudy (australopithecus):

anything you are not getting?? solve for x and add the number of visitors from both months

OpenStudy (anonymous):

would it be 1930?

OpenStudy (australopithecus):

x = 1930 since, "x represents the number of visitors to the site the previous month." then we know the amount of visitors for the previous month

OpenStudy (australopithecus):

now it is just the matter of adding two numbers together

OpenStudy (anonymous):

x-0.3x=1351=> 0.7x=1351=> x=1930 The total visitors of the 2 months =1930+1351 =3,281

OpenStudy (anonymous):

Hit the best answer button

OpenStudy (anonymous):

okay